OpenStudy (anonymous):

x/2-3/5=3/2 ~ ok, I understand that I have to start off by determining the LCD, but I'm lost after that. I need to solve for x.

OpenStudy (anonymous):

what's your LCD?

OpenStudy (anonymous):

I believe it's 10

OpenStudy (anonymous):

yes it is, now multiply this by entire equation like this: \[10(x/2) - 10(3/5) = 10*3/2\]

OpenStudy (anonymous):

the equation reduces to this: 5x-6=15, can you proceed from here?

OpenStudy (anonymous):

your gonna hate me ~ but I'm still lost on the first stuff you posted. I'm 49 years old & this is my first ever algebra. I'm sorry

OpenStudy (anonymous):

OK, from first part of your equation, there is an LCD of 10 multiplied by x/2, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't understand how to multiply x/2

OpenStudy (anonymous):

 well you just cancel by the 2 and it reduces as such